OpenStudy (anonymous):

can someone please explain EACH step that are necessary to Solve 4x-2(x+3)=x+1 , like explain how you get the answer.

OpenStudy (anonymous):

4x-2(x+3)=x+1 4x-2x-6=x+1 2x-6=x+1 x-6=1 x=7

OpenStudy (anonymous):

can you explain how you got that please

OpenStudy (anonymous):

4x-2(x+3)=x+1 (given equation) 4x-2x-6=x+1 (distribute the -2 over the binomial) 2x-6=x+1 (combine the 4x and 2x) x-6=1 (subtract x from both sides) x=7 (add 6 to both sides)